OpenStudy (anonymous):

Solve the equation 6n + 11 = -7 + 3n for n. n equals negative two n equals negative six n equals six n equals two im giving medals!!!!!!!!

OpenStudy (luigi0210):

Moves variables to one side and numbers to the other..: \[\LARGE 6n-3n=-7-11\] \[\LARGE 3n=-18\] Finish it